# Conditional Latent Space Molecular Scaffold Optimization for Accelerated Molecular Design

### v1

@@ 5.2 QED Optimization
5.2.1 Model Benchmark
Table 2 shows the top QED values achieved by various models, highlighting CLaSMO’s performance. Our method reaches a top QED score of 0.948, matching state-of-the-art results from PGFS Gottipati et al. (2020) and GP-MOLFORMER Ross et al. (2024). Among these models, Scaffold-GGM, DrugHIVE, and PGFS are substructure-addition-based methodologies, while the others generate molecules from scratch. A key distinction is that CLaSMO’s CVAE is trained only on substructures, not complete molecules, using a dataset of only 18,706 instances. The maximum QED value of the substructure among these instances is 0.53, with average QED value overall 0.427333Distribution of the QED values used in CVAE training is provided in the Appendix A.2.. In contrast, competitor models are trained on significantly larger datasets and full molecules, often encountering molecules with high QED values during training. These results underscore CLaSMO’s efficiency and lower computational cost, demonstrating the advantages of using carefully curated datasets and tailored methodologies for the target task.
Table 2: Comparison of top QED values achieved by various models with their dataset sizes.
Model Top QED Dataset Size
JT-VAE Jin et al. (2018) 0.925 250,000
Scaffold-GGM Lim et al. (2020) [^5] Scaffold-GGM results obtained by running their released code for QED task. 0.928 349,809
DrugHIVE Weller & Rohs (2024) 0.940 27 Million
LIMO Eckmann et al. (2022) 0.947 250,000
PGFS Gottipati et al. (2020) 0.948 -
GP-MOLFORMER Ross et al. (2024) 0.948 1.1 Billion
CLaSMO (ours) 0.948 18,706
5.3 Docking Simulation Score Optimization
In this section, we present our docking simulation score optimization results using the KAT1 protein as the target. KAT1 is an ion channel protein, and compounds with favorable docking scores are more likely to bind effectively and influence its function, which is crucial for developing therapeutic agents targeting ion channel-related conditions [^4] Details of the KAT1 protein can be found at https://pdbj.org/mine/summary/6v1x.. By optimizing docking scores, we aim to identify compounds with the potential to modulate ion channel function and develop novel therapeutic strategies. Unlike the relatively simple calculation of QED values, optimizing docking scores for KAT1 requires computationally intensive simulations. For these calculations, we used Schrödinger Software Schrödinger (2023), known for providing reliable and accurate results, with each simulation taking anywhere from a few minutes to several hours. Given the high computational cost of these simulations, CLaSMO’s sample-efficient approach is particularly valuable, enabling effective optimization while minimizing the number of expensive evaluations.
Similar to the QED optimization setting, the search domain for each latent dimension in $\mathcal{Z}$ was set to $[-6,6]$. To address the high cost of obtaining labeled data for docking scores, we leveraged the same training set used in the QED experiments, treating them as low-fidelity instances for training the GP model. This approach allows the surrogate model to learn the relationships between atoms and latent vectors, even though the labels originate from QED rather than docking scores. By reusing this shared data, we enhance the GP model’s capacity to generalize and effectively guide the optimization process, reducing the dependency on extensive docking score evaluations. Given the significant time required for each docking score calculation, we limited the experiments to 10 compounds from the ZINC250K dataset and ran CLaSMO for 100 iterations per compound using two Dice similarity thresholds, $\tau\in[0.25,0.50]$. This setup ensures that despite the computational expense, we can still obtain comprehensive and meaningful results for docking score optimization, and observe if CLaSMO is capable of optimizing docking scores under similarity constraints.
We present the distribution of docking scores obtained from both CLaSMO experiments in Fig. 7. The results clearly demonstrate that CLaSMO, at both similarity thresholds, achieves significant improvements in docking scores compared to the initial scaffolds. Figure 7 provides detailed metrics, including the maximum and average improvements across the models, as well as the initial best values. Notably, CLaSMO with a similarity threshold of $\tau=0.25$ achieved improvements of up to 96.3% over the initial scaffold docking score, while CLaSMO with $\tau=0.50$ achieved improvements of up to 75.1%, both indicating strong optimization performance despite the limited number of LSBO iterations, showcasing CLasMO’s sample-efficiency. In Fig. 8, we provide example molecules obtained from our experiment, where we observe that, even at the beginning stages of the optimization where $\text{DICE}(\bm{S},\bm{S^{\prime}})$ still above 0.70, we observe substantial improvements.
Figure 6: Distribution of docking scores for CLaSMO optimization experiments. Lower docking scores indicate better binding affinity, highlighting the efficacy of CLaSMO in enhancing docking performance across different similarity constraints.
     Metric Input $\bm{\tau=0.25}$ $\bm{\tau=0.50}$ Min DS -5.051 -9.30 -8.57 Mean DS -4.934 -8.32 -6.62 Mean Imp. - 70.1% 35.4% Max Imp. - 96.3% 75.1% Figure 7: Summary of docking score (DS) optimization results: Max and mean DS values from input scaffolds and CLaSMO experiments with various $\tau$ levels, and mean and max improvement rates over DS values of input scaffold.
Figure 8: Examples of molecules obtained from CLaSMO $\tau=0.25$ experiments from docking score optimization setting. In both rows, molecules in the left refer to the input scaffolds. Each molecule to the right shows one step of substructure addition with their resulting docking scores. Docking scores are significantly improved by adding small substructures to different regions of the input molecule.
@@ 5.4 Ablation Study
We conducted an ablation study to evaluate the impact of incorporating atomic environment features as condition vectors in the generative process. In the training of the baseline VAE model, only substructures of the molecules were used, without conditioning on atomic properties. This setup allowed us to isolate the effect of conditioning on scaffold optimization.
We repeated the QED optimization experiments using the VAE model. The QED optimization results, shown in Table 3, demonstrate that incorporating atomic environment features in the CVAE significantly improves the generation of substructures. The conditioning mechanism enhances the model’s ability to generate substructures that bond more effectively with the scaffold, leading to higher success rates in optimizing molecular properties. While the VAE without conditioning still gains from the LSBO framework’s efficient optimization, incorporating atomic properties as conditions significantly enhances the quality of the generated molecules.
Table 3: QED optimization results: Comparison of VAE and CVAE at different $\tau$ thresholds, reporting mean and max improvement rates.
Metric $\bm{\tau=0.25}$ $\bm{\tau=0.50}$
Mean Imp. (VAE) 12.31% 8.41%
Mean Imp. (CVAE) 19.14% 16.08%
Max Imp. (VAE) 68.16% 32.28%
Max Imp. (CVAE) 114.57% 74.74%
@@ 5.5 Interactive Optimization
This interactive approach offers several key advantages. It enables the integration of expert knowledge into the optimization process, allowing CLaSMO to operate in a Human-in-the-loop setting. Experts can leverage their domain-specific insights to target specific regions they find promising, ensuring that modifications are not only data-driven but also aligned with scientific understanding. Meanwhile, CLaSMO maintains its efficient optimization process, using LSBO to enhance molecular properties and preserve sample efficiency. Detailed user guidelines for this open-source web application are provided in the Appendix A.
@@ 6 Conclusion
In this paper, we introduced CLaSMO, a novel framework that combines CVAE and LSBO for scaffold-based molecular optimization. Our approach efficiently explores latent spaces to optimize molecular properties, such as QED and docking scores, while maintaining structural similarity with the input scaffold to improve the chances of real-world viability of optimized molecules. By conditioning substructure generation on the atomic environment of the target region in the input molecule, CLaSMO generates chemically meaningful modifications. The experimental results demonstrate that CLaSMO achieves state-of-the-art performance, even with limited training data, a significantly smaller model compared to existing methods, and a limited number of optimization iterations, showcasing its sample-efficiency and lower computational cost. Furthermore, CLaSMO’s ability to control structural divergence through similarity constraints ensures robust performance across different optimization tasks. Although this paper focuses on scaffold-based modifications, CLaSMO is fully compatible with whole molecules, requiring no changes to its methodology. Additionally, we have open-sourced a web application to allow chemical experts to use CLaSMO in a Human-in-the-Loop setting, further extending its practical applicability. Overall, CLaSMO exemplifies the power of combining scaffold-based strategies with LSBO, offering a highly effective tool for targeted drug discovery and broader molecular design challenges.
@@ A.1 Definitions of Atomic Features
A.2 QED Value Distribution of Input Substructures
Figure 9 shows the distribution of QED values for the substructures used to train the CVAE model, generated through our data preparation process.
Figure 9: QED value distribution of substructures derived from our data preparation methodology.

### v2

@@ 5.2.2 Property Optimization Results
In summary, even without similarity constraints, CLaSMO effectively explores the chemical space to optimize molecular properties with high sample-efficiency. In this setting, MARS and Graph-GA methodologies provides more competitive results compared to constrained optimization settings. However, CLaSMO still achieves the best performance across 11 out of 20 optimization tasks [^8] In CLaSMO, the ratio of generated substructures that successfully bond with the given scaffold serves as an important metric for assessing optimization performance. Appendix A.9 provides a detailed analysis of bonding success rates throughout the optimization process.
Across all experimental settings, we observe notable differences in sample-efficiency among the benchmarked methods. MolDQN, a reinforcement learning based approach, and SynFlowNet, a GFlowNet-based approach consistently require substantially more oracle evaluations to achieve comparable optimization levels. These findings are consistent with those reported by (Gao et al., 2022), which highlighted the low property optimization capabilities of such methods [^9] SynFlowNet, unlike MolDQN, was not included in the benchmark conducted by Gao et al. (2022). However, other GFlowNet-based methodologies evaluated in that study exhibited lower optimization performance compared to alternative approaches.. Our results further reinforce that their slower convergence makes them less suitable for resource-constrained sample-efficient optimization tasks, where a low evaluation budget is available. In contrast, genetic algorithm-based methods, especially Graph-GA, remain competitive and often perform well across a variety of tasks. MARS, on the other hand, also shows strong performance, particularly in multi-property optimization scenarios, and achieves solid results overall [^10] Further discussion on the performance comparison on MARS and CLaSMO is provided in Appendix A.10. However, CLaSMO consistently outperforms all baselines across the majority of tasks and similarity constraint levels. These results highlight CLaSMO’s robustness and efficiency across both constrained and unconstrained optimization settings, where sample-efficiency is essential due to limited evaluation budgets.
To illustrate how CLaSMO optimizes input scaffolds, Fig. 8 presents two representative examples from the QED optimization task. In each case, CLaSMO incrementally modifies the input scaffold by adding substructures that are conditionally generated based on the local atomic environment. Each step in the optimization process yields a molecule with an improved QED score, demonstrating the model’s ability to progressively refine structure–property relationships. The examples show QED improvements from 0.615 to 0.938 and from 0.827 to 0.948, underscoring CLaSMO’s ability to produce chemically valid and property-enhancing modifications through its conditional generation mechanism and LSBO-guided search.
5.2.3 Synthetic and Retrosynthetic Accessibility Evaluation
To assess the synthesizability of the optimized molecules, we analyzed both their synthetic accessibility (SA) scores and retrosynthetic accessibility (RA) scores.
The SA score estimates how difficult a compound is to synthesize based on structural complexity and fragment contributions. Lower SA scores indicate molecules that are easier to synthesize, with values typically ranging from 1 (very easy) to 10 (very difficult). Scores above 5 are generally considered indicative of challenging synthesis Anede et al. (2024).
Given that CLaSMO and MARS exhibited the strongest performance in property optimization, we focused our SA and RA analysis on these two methods [^11] There is a strong correlation between the SA score and the similarity to the original molecule. When the similarity to the original molecule is high, the SA score also tends to be favorable, as the original molecule is synthesizable. Conversely, when the similarity is low, the SA score tends to be unfavorable. As a result, molecules generated by methods that involve minimal modifications to the original molecule tend to receive good SA scores. Therefore, we focus on comparing CLaSMO and MARS, both of which achieve high scores in property optimization.. We compared the average SA scores obtained under the lowest and highest similarity constraints, $\tau\in\{0,0.50\}$. Figure 9 presents the distributions of these scores. The horizontal dashed line represents the average SA score of the input scaffolds, which is 3.00.
Both methods generate molecules that are, on average, more complex than the input scaffolds, as the optimized molecules exhibit higher molecular weight and structural complexity. While such increases can pose challenges for synthesis, our results demonstrate that with targeted, small-scale modifications, good synthesizability can still be maintained. At $\tau=0$, the average SA score is 3.88 for CLaSMO and 3.74 for MARS. At $\tau=0.50$, these scores improve slightly to 3.73 and 3.68, respectively. Although MARS consistently produces molecules with slightly better synthetic accessibility, CLaSMO maintains competitive synthesizability while achieving stronger performance in property optimization. As indicated by the distribution of SA scores, the generated molecules consistently score below 5, suggesting that their synthesis is likely to be straightforward (Anede et al., 2024).
To complement this structure-based evaluation, we further report the RA score, which offers a model-driven estimate of a compound’s synthesizability based on its likelihood of being successfully deconstructed by retrosynthetic planning algorithms. The RA score is computed using a machine learning classifier trained on retrosynthesis outcomes and ranges from 0 to 1, with higher values indicating a higher predicted probability that a valid synthetic route exists. In our evaluation, we report the percentage of molecules that are classified as synthesizable.
As shown in Figure 10, 100% of the input scaffolds are deemed synthesizable. Among the optimized molecules from the experiments with similarity threshold $\tau=0.5$, 88.0% of those generated by CLaSMO and 90.1% of those from MARS are classified as synthesizable. For the case where there is no similarity threshold, $\tau=0$, 79.2% of those generated by CLaSMO and 80.6% of those from MARS are classified as synthesizable. These results support the notion that both methods yield molecules that are not only structurally accessible (as evidenced by favorable SA scores) but also recognized by data-driven retrosynthetic models as viable targets for synthesis.
Together, the SA and RA evaluations provide a robust and complementary view of synthesizability. While SA scores reflect fragment-level and structural complexity, RA scores offer a practical assessment from the perspective of route planning. CLaSMO, in particular, achieves a compelling balance, delivering high amount of property improvements across many tasks while maintaining strong performance across both accessibility metrics and similarity constraints. These findings reinforce its potential for real-world molecular design scenarios where synthetic feasibility is essential.
Figure 9: Distributions of average SA scores for molecules optimized by CLaSMO and MARS under similarity constraints $\tau=0$ (A) and $\tau=0.50$ (B). Lower scores correspond to higher synthesizability.
Figure 10: RA scores for input scaffolds and molecules generated by CLaSMO and MARS. Bars indicate the percentage of molecules predicted to be synthesizable by a retrosynthesis planning model.
@@ 5.3 Ablation Study
To assess the individual contributions of each component in the CLaSMO framework, we conducted ablation studies on (i) the role of LSBO as the search strategy and (ii) the effect of conditioning on atomic environment features in the generative model. All evaluations were conducted in a 2D latent space for consistency and under a high similarity constraint ($\tau=0.5$).
In the first setting, we use a standard VAE without conditioning, paired with LSBO, denoted as LSBO+VAE. In the second, we use the full CVAE model but replace LSBO with random search (RS), denoted as RS+CVAE. These are compared against the original LSBO+CVAE configuration used in CLaSMO. To evaluate these alternatives, we report the final Top-10 average scores across three representative oracle optimization tasks: mestranol_similarity, QED, and troglitazone_rediscovery.
As shown in Table 1, LSBO+CVAE achieves the highest performance, highlighting the synergy between guided search and conditional generation. LSBO+VAE underperforms compared the other methods, indicating that lack of conditioning leads to chemically incompatible fragments that are penalized under high similarity constraints. While LSBO provides sample-efficient search, it cannot overcome this semantic mismatch. RS+CVAE outperforms LSBO+VAE in two out of three tasks, demonstrating that conditioning helps narrow the search space to chemically meaningful and similarity-compliant substructures—even when sampled randomly. These results further underscores the importance of scaffold-aware conditioning in constrained optimization settings.
Together, these findings confirm that both components, conditional generation and LSBO, are essential for achieving high performance.
Table 1: QED optimization results from the ablation study. We compare three configurations: LSBO+CVAE (original setting), LSBO+VAE (without conditioning), and RS+CVAE (without guided optimization). Results are reported as Top-10 average QED scores under a similarity constraint of $\tau=0.5$. The full CLaSMO configuration achieves the highest performance.
Oracle LSBO+CVAE RS+CVAE LSBO+VAE
mestranol_similarity 0.331 0.303 0.308
QED 0.930 0.922 0.897
troglitazone_rediscovery 0.239 0.216 0.191
@@ 5.4 Interactive Optimization
This interactive approach offers several key advantages. It enables the integration of expert knowledge into the optimization process, allowing CLaSMO to operate in a Human-in-the-loop setting. Experts can leverage their domain-specific insights to target specific regions they find promising, ensuring that modifications are not only data-driven but also aligned with scientific understanding. Meanwhile, CLaSMO maintains its efficient optimization process, using LSBO to enhance molecular properties and preserve sample-efficiency. Detailed user guidelines for this open-source web application are provided in the Appendix A.12.
@@ 6 Conclusion
In this paper, we introduced CLaSMO, a novel framework that combines CVAE and LSBO for scaffold-based molecular optimization. Our approach efficiently explores latent spaces to optimize molecular properties in wide variety of objectives, while maintaining structural similarity with the input scaffold to improve the chances of real-world viability of optimized molecules. By conditioning substructure generation on the atomic environment of the target region in the input molecule, CLaSMO generates chemically meaningful modifications. The experimental results—encompassing tasks such as rediscovery of molecules and multi-property optimization settings—demonstrate that CLaSMO consistently delivers superior performance across diverse settings and optimization objectives of varying complexity. By utilizing a pioneering LSBO-based molecular editing framework that exhibits superior sample-efficiency, CLaSMO proves to be a highly effective methodology for low-budget molecular optimization scenarios. Furthermore, CLaSMO’s ability to control structural divergence via similarity constraints enables consistent performance across diverse optimization tasks, while also yielding favorable outcomes in terms of synthetic accessibility. Although this paper focuses on scaffold-based modifications, CLaSMO is fully compatible with whole molecules, requiring no changes to its methodology. Additionally, we have open-sourced a web application to allow chemical experts to use CLaSMO in a Human-in-the-Loop setting, further extending its practical applicability. Overall, CLaSMO exemplifies the power of combining scaffold-based strategies with LSBO, offering a highly effective tool for targeted drug discovery and broader molecular design challenges.
Limitations and Scope
As discussed throughout this manuscript, CLaSMO is particularly well-suited for low-budget optimization scenarios where sample-efficiency is critical. It is inherently designed to perform additive modifications on a given scaffold, which gradually increases the molecular size and weight with each editing step. By operating on scaffolds, CLaSMO helps limit the extent of molecular growth, ensuring that the optimized molecules do not become too large compared to their original counterparts, with the help of its constrained optimization capabilities. However, the current framework does not support the removal or replacement of existing substructures, limiting its flexibility in certain optimization settings. CLaSMO may not be ideal in scenarios where maintaining a fixed molecular size is crucial, or where more flexible transformations—such as fragment replacement or crossovers between molecular series—are required. In such settings, its utility may be limited without integration into broader pipelines. Besides, with its emphasis on introducing small modifications to the input scaffolds and given the scope of this study, CLaSMO uses a set of small fragments in its training dataset. Depending on the requirements of the problem, training on more complex fragments may be necessary, as discussed in Appendix A.2.
Another limitation, related to the discussion above, is that CLaSMO is not suitable for de novo, from-scratch molecule generation. While it is effective at adding substructures to a given molecule, it is not designed to construct molecules from scratch. In principle, CLaSMO could be adapted to grow a molecule starting from a single atom or fragment by iteratively generating substructures, but such usage falls outside the intended design and is beyond the scope of this work.
Nonetheless, CLaSMO remains a valuable tool when scaffold-based additions are permissible. It can be effectively combined with other methods presented in this study, serving as a component within a modular molecular optimization framework tailored to diverse design constraints.
@@ A.1 Definitions of Atomic Features
A.2 Training with BRICS Fragments from the ZINC250K Dataset
Our initial decision to use BRICS fragments derived from the QM9 dataset was motivated by our goal of training a generative model capable of making small, chemically interpretable modifications to input scaffolds. However, the CLaSMO framework is not limited to the use of QM9. Depending on the task at hand, the CVAE can be trained on BRICS fragments from alternative datasets that offer different levels of structural complexity and diversity.
As discussed in Section 5.1.1, BRICS fragments obtained from ZINC250K exhibit a wider range of sizes and chemical diversity compared to those from QM9. In this section, through simple experiments, we investigate the implications of using the ZINC250K dataset within the CLaSMO framework.
A.2.1 Training CVAE with Only ZINC250K Fragments
We first conduct a preliminary study in which we train the CVAE model using the complete set of BRICS fragments extracted from ZINC250K, without applying any molecular weight filtering. This results in 35,232 unique BRICS fragments and their corresponding atomic environment features, significantly increasing the chemical diversity and structural complexity available during training.
To assess the effect of this increased diversity and fragment size, we trained CVAE models using this ZINC250K-derived dataset with latent dimensions of 2 and 8, keeping the model architecture, hyperparameters, and training procedure consistent with the QM9-based setting. The model with a 2-dimensional latent space achieved a test reconstruction accuracy of 87.1%, while the model with an 8-dimensional latent space achieved 91.2%. Despite the added complexity of the ZINC250K fragments, the 2D model still performed reasonably well, though further improvements could likely be achieved with proper hyperparameter tuning.
We then evaluated the property optimization performance of the 2D ZINC250K-trained model in a QED optimization task and benchmarked it against the original CLaSMO model trained on QM9-derived fragments. We also compared the average molecular weight of the optimized molecules.
As shown in Table 3, the use of ZINC250K-derived fragments results in the generation of molecules with substantially higher molecular weights—averaging 473.12 Da—due to the inherently larger size of the training fragments. Under the $\tau=0$ similarity setting, the ZINC250K-trained model achieves a top-10 average QED of 0.890. This performance surpasses all baseline methods except MARS, which it closely approaches, demonstrating that the model remains effective even when operating with a more structurally complex fragment vocabulary. In the following subsection (Appendix A.2.2), we show that using small fragments filtered from the ZINC250K dataset in combination with the QM9 dataset is more suitable for our model, which consistently outperforms the baselines, including MARS.
However, under the more restrictive $\tau=0.5$ setting, the model’s performance drops more noticeably, achieving a QED of only 0.871. This decline can be attributed to the larger substructures generated by the model, which make it difficult to maintain high similarity to the input scaffold. The model struggles to produce small, localized modifications, which are essential under stricter similarity constraints. In contrast, the original QM9-based CLaSMO setting maintains high performance across both similarity thresholds, highlighting the importance of fragment size in preserving molecular similarity during optimization.
It is also important to note that the current model architecture was originally designed for QM9-sized fragments and has not been adapted to handle the increased complexity introduced by the ZINC250K fragments, many of which exceed 400 Da in molecular weight. With a more carefully selected model architecture and appropriate hyperparameter tuning—such as latent dimensionality, condition embedding size, and decoder capacity—we expect this setup to yield stronger performance. In the next substructure, we show a more suitable alternative to scale CLaSMO with new fragments.
Table 3: Comparison of average optimized molecular weight (Avg. Mol. Weight) and top-10 average QED scores for models trained on BRICS fragments from different datasets.
CLaSMO Setting
Avg. Mol. Weight
$\bm{\tau=0}$
QED
$\bm{\tau=0}$
QED
$\bm{\tau=0.50}$
ZINC250K Fragments, LD:2 473.12 0.890 0.871
QM9 Fragments, LD:2 (Original Setting) 320.08 0.932 0.930
MARS 313.28 0.896 0.890
A.2.2 Training CVAE with a Mixture of ZINC250K and QM9 Fragments
As discussed in Section 5.1.1, the ZINC250K dataset contains some BRICS fragments that are comparable in size to those extracted from QM9. By filtering for such fragments, we curated a subset of 4,706 unique BRICS fragments along with their corresponding atomic environment vectors. This subset enhances the chemical diversity of the training data while maintaining the small fragment size necessary for fine-grained scaffold editing.
We augmented the QM9-derived dataset with a filtered subset of BRICS fragments from ZINC250K and trained a CVAE model using a 2-dimensional latent space, consistent with the original setup. No additional hyperparameter tuning was performed; all settings from the original CLaSMO configuration were reused for both model training and the LSBO procedure.
In the QED optimization task, the mixed-fragment model achieved a Top-10 average QED of 0.912—surpassing all baseline models and closely approaching the performance of the original CLaSMO setup. The average molecular weight of the generated molecules was 327.04, remaining aligned with the QM9-only setting. We further evaluated this configuration on the troglitazone_rediscovery and mestranol_similarity benchmarks with no similarity constraint ($\tau=0$). As shown in Table 4, the model outperformed the strongest baseline method, MARS, and performed comparably to the original CLaSMO setting across all metrics, confirming the method’s robustness to variation in fragment sources.
These results highlight CLaSMO’s strong generalization capabilities, even when exposed to more structurally diverse training data. Notably, the model achieved this level of performance without any additional tuning, suggesting that task-specific or dataset-aware hyperparameter optimization could lead to further improvements. This demonstrates CLaSMO’s potential to scale effectively with alternative datasets composed of small molecular fragments.
Table 4: Comparison of CLaSMO performance across three tasks—QED optimization, troglitazone_rediscovery, and mestranol_similarity—under different fragment settings and similarity constraint $\tau=0$. The setting that combines QM9 and ZINC250K fragments demonstrates comparable performance to the original QM9-only configuration and outperforms the MARS baseline across all tasks, highlighting CLaSMO’s robustness and potential for scaling with diverse fragment sources.
CLaSMO Setting QED troglitazone_rediscovery mestranol_similarity
QM9 + ZINC250K Fragments, LD:2 0.912 0.233 0.342
QM9 Fragments, LD:2 (Original Setting) 0.932 0.243 0.355
MARS 0.896 0.216 0.286